A retail store can sell 20 hard disk drives per week at a price of $400 each. The manager estimates that for each $10 reduction in price, she can sell two more drives each week.
a. Express the revenue function in terms of selling price.
b. What price will maximize revenue?

Answer by ewatrrr(24785)

 
Hi,
A retail store can sell 20 hard disk drives per week at a price of $400 each. The manager estimates that for each $10 reduction in price,
she can sell two more drives each week.
a. Express the revenue function in terms of selling price. R=(20+2x)($400-10x)
b. What price will maximize revenue? x = 15 $400 - 150 = $250
R=(20+2x)($400-10x) = - 20x^2 + 600x + 8000 = -20(x-15)^2 + 4500 + 8000
= -20(x-15)^2 + $12,500